OpenStudy (anonymous):

What is the area of the square in square meters and square centimeters? A. 6 m² and 600 cm² B. 36 m² and 3600 cm² C. 36 m² and 36,000 cm² D. 36 m² and 360,000 cm²

OpenStudy (anonymous):

Is there a diagram/picture?

OpenStudy (anonymous):

Area of square = a^2

OpenStudy (anonymous):

A

OpenStudy (anonymous):

no, you have to square it

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

6^4=1,296

OpenStudy (anonymous):

no..base* base

OpenStudy (anonymous):

6*6=?

OpenStudy (anonymous):

6*6

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

36

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Yes its 36

OpenStudy (anonymous):

convert is to cm^2 cm now, you have 36m^2, how many meters in a cm?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

3600

OpenStudy (anonymous):

Hmm.actually 1m^2 = 10,000cm^2 So 10,000 * 36

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

360,000

OpenStudy (anonymous):

There you go

OpenStudy (anonymous):

Yes, so our answer is 36m^2 or 360,000cm^2

OpenStudy (anonymous):

I wanted you to convert from cm to cm^2 as it's a good habit and good to understand, so please I suggest you learn how to convert, as it's very essential in higher level courses!